

********************************************************************
UARS WEEKLY STATUS REPORT - 11 Jan 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods. We supported a telecon with the FOT, SUSIM and
other project personnel to discuss the move to SSPP closed loop solar
tracking.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry. On Jan. 10th, SUSIM
participated in a meeting at GSFC concerning the use of SSPP
Closed Loop Mode.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM2K:: node was marked disabled on the CDHF and PEM is having
some difficulty finding someone to reset the PEM2K transfer status. This
line should be marked as enabled.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 December 2001 (Orbit 56312, GMT Day 363)
through 04 January 2002 (Orbit 56417, GMT Day 004).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere
viewing) and is performing nominally under two-battery
operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
---------- ---- ----- -------
HALOE 363/1414 56321 Power ON (START)
HALOE 364/0217 56329 Power ON - COMPLETE
WINDII 002/1055 56379 Power ON
PEM AXIS 004/0039 56403 POWER ON (7 WATTS)
HALOE 006/1748 56443 Power OFF
The beta angle ranged between the angles of 30.94 degrees to
09.68 degrees for this report period. The beta angle is now
decreasing towards a minimum angle of 0 degrees on 03 February
2002 (DOY 034).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3
are currently maintaining greater than 26.4 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a
temperature delta between Battery 2 and Battery 3 of 0.97 to 1.31
degrees C. Additional battery performance data for each battery
on 22 December 2001 and 28 December 2001 are provided in
Appendix D.
The clock error ranged from -8.9 to 8.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the
initial excursion of ALERTS or Out-Of-Limits conditions detected
by the FOT for this report. A summary is listed below and the
details of ALERT and Out-Of-Limits occurrences are listed in
Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the
listed operational element(s) during this report period (see
Appendix A).
FOT = 3
PEM/AXIS = 1
HALOE = 3
WINDII = 1
UARS Anomaly Reports initiated during this report period are
listed in Appendix E and summarized below.
FOT = 0
Zero (0) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
* Transponder A: 01/333 - 2287.496855
* Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period.
Anomaly Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued its testing with the new TAC systems this week.
The FOT in conjunction with the TAC developers are still trying
hard to resolve the open issues on the TAC system. Several more
tests were run in an attempt to isolate the problem without
success. Only a few tests remain that could give any further
insight into the problem.
ECLIPSE procedure and page validation will be the focal point of
the reengineering effort in the upcoming weeks. The simulator
will be used to test the procedures and real time telemetry will
be used in side by side comparison to validate pages. A realtime
test will be conducted on 09 January 2002 where a NOOP command
will be sent from the Eclipse system to the spacecraft.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday,
January 03, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 03 February 2002 (DOY
034). This maneuver will be a Reverse to Forward Flight maneuver
(Southern Hemisphere viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 January 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56321 363/1414 01-745 HALOE Power ON (START)
56327 364/0037 01-747 NBTR NOOPS
56329 364/0217 01-746 HALOE Power ON - COMPLETE
56379 002/1055 02-001 WINDII Power ON
56403 004/0039 02-002 PEM AXIS POWER ON (7 WATTS)
56417 005/0254 02-003 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
56443 006/1748 02-004 HALOE Power OFF
56444 006/1929 02-005 OBC Clk Adj fr 20.6 to 0.00 msec/day
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 January 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 January 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
56417 005/0300 Clk Adj 0.00 to 20.6 msec/day
56444 006/1929 Clk Adj fr 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 January 2002
29 December 2001 - GMT Day 363 Beta = 14.96 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ----- -----
1 -560.0/-728.0 -4.08/-4.99 +0.0/ -0.0 10.72 0.000 0.00
2 +000.0/-224.0 +4.15/+3.50 +28.8/-12.8 26.40 1.020 19.46
3 +224.0/-112.0 +2.85/+2.52 +27.2/-12.8 26.24 1.020 18.99
04 January 2002 - GMT Day 004 Beta = 40.87 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ----- -----
1 -560.0/-728.0 -3.77/-4.69 +0.0/ -0.0 10.72 0.000 0.00
2 +112.0/-952.0 +5.14/+4.48 +26.8/-14.4 25.76 1.030 20.51
3 +224.0/-112.0 +4.48/+3.17 +26.4/-14.4 25.76 1.020 20.39
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 January 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-75 FOT see list Dec 01 14170 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
SEVNTFW02/01